Computer Company assembles personal computers and sells them in the retail marketplace. The company is organized into two profit centers: the assembly division and the distribution division. The demand curve facing the company (and the distribution division) is P=3,500 – 10Q. The marginal cost for assembly (which includes purchasing the parts) is constant at $450. The distribution division faces constant marginal distribution costs of $50 per unit.

A. What is the profit-maximizing retail price and output for the firm as a whole?

B. If the assembly division has monopoly power to set the transfer price, what transfer price will it select (assuming it knows all the information above)? Calculate the profits for the two divisions in this case.


1
Expert's answer
2021-12-07T13:09:20-0500

A. The profit-maximizing retail price and output for the firm as a whole are:

MR = MC,

MC = 450 + 50 = 500,

"MR = TR'(Q) = 3,500 - 20Q,"

3,500 - 20Q = 500,

Q = 200 units,

"P = 3,500 - 10\u00d7200 = 1,500."

B. If the assembly division has monopoly power to set the transfer price, then the transfer price can be not less then the total marginal cost, which is 500.